# Fast Generation of Feasible Trajectories in Direct Optimal Control

### v1

@@ Abstract
This paper examines the question of finding feasible points to discrete-time optimal control problems. The optimization problem of finding a feasible trajectory is transcribed to an unconstrained optimal control problem. An efficient algorithm, called FP-DDP, is proposed that solves the resulting problem using Differential Dynamic Programming preserving feasibility with respect to the system dynamics in every iteration. Notably, FP-DDP admits global and rapid local convergence properties induced by a combination of a Levenberg-Marquardt method and an Armijo-type line search. The efficiency of FP-DDP is demonstrated against established methods such as Direct Multiple Shooting, Direct Single Shooting, and state-of-the-art solvers.
I INTRODUCTION
@@
Reducing infeasibility and finding feasible points is a key building block of constrained optimization algorithms. State-of-the-art nonlinear optimization solvers such as IPOPT [21] or filterSQP [9] employ a feasibility restoration phase to avoid cycling, to overcome infeasible subproblems and failure of sufficient decrease by iterating towards the feasible set. Feasible iterate solvers such as SEQUOIA [16] or FSQP [24] require every iterate, including the initial guess, to be feasible. Since giving feasible initial guesses for a generic NLP can be difficult, an initial feasibility phase solves a feasibility problem to find a feasible initial guess. The feasibility-projected Sequential Quadratic Programming (SQP) framework [18], [22] is a feasible iterate method that exploits structure to solve linearly constrained Optimal Control Problems (OCPs). Their method employs a trust-region SQP method that uses a feasibility projection to achieve feasibility in every iterate. The feasibility projection is based on Differential Dynamic Programming (DDP) to guarantee feasibility with respect to the dynamics constraints while solving an additional optimization problem to maintain feasibility with respect to linear constraints. Similarly, the method proposed in this paper is based on DDP but relies on reformulating additional inequality constraints in terms of quadratic penalties in the objective function.
@@
DDP was initially introduced by Mayne in [15]. It is a method to solve unconstrained OCPs that performs a Riccati backward recursion for a quadratic approximation of the unconstrained OCP and a nonlinear forward simulation to achieve feasibility of the dynamics in every iteration. Local quadratic convergence for exact Hessian DDP was first shown in [6]. The first global convergence proof for line search DDP was presented in [23]. In [2], it was shown that DDP, Direct Multiple Shooting (DMS), and Direct Single Shooting (DSS) recover the same local convergence rates when used with a Generalized Gauss-Newton (GGN) Hessian approximation. Constrained optimization algorithms require a globalization mechanism, where both the decrease of the objective function and the infeasibility of the constraints are controlled. Since the relative weighting of these two terms is a nontrivial task, DDP is used in this paper as a means to avoid globalization of the dynamics constraints.
@@
The contributions of this paper are threefold. Firstly, the optimization problem of finding a feasible trajectory is discussed starting from state-of-the-art solvers and a suitable structure-preserving feasibility problem is identified for the special case of OCPs. Secondly, based on the proposed optimization problem, a globally convergent DDP algorithm is introduced. Finally, the novel feasibility algorithm is compared against state-of-the-art methods, DMS, and DSS on challenging OCPs.
@@
The remainder of this paper is structured as follows. Section II discusses the optimization problem of finding a feasible point. In Section III, the novel feasibility algorithm is introduced. Simulation results are presented in Section V. Section VI concludes this paper.
@@
The real numbers are denoted by $\mathbb{R}$. The concatenation of column vectors $w_{1}\in\mathbb{R}^{n_{1}}$ to $w_{n}\in\mathbb{R}^{n_{n}}$ is denoted by $(w_{1},\ldots,w_{n})\in\mathbb{R}^{n_{1}+\ldots+n_{n}}$. Iteration indices are denoted by a letter $k$ and superscripts in parentheses, e.g., $w^{(k)}\in\mathbb{R}^{n_{w}}$. For a given iterate $w^{(k)}$, quantities evaluated at that iterate will be denoted with the same superscript, e.g., $f^{(k)}$ or $\nabla f^{(k)}$. In the context of optimal control problems, an OCP stage is denoted by a letter $i$ and subscript, e.g., $x_{i}$ and $u_{i}$. Let $[\cdot]^{+}\colon\mathbb{R}^{n}\to\mathbb{R}^{n},x\mapsto\max\{0,x\}$ denote the componentwise maximum of a vector with the zero vector. Finally, $f(w)=\mathcal{O}(g(w))$ for functions $f,g\colon\mathbb{R}^{n_{w}}\to\mathbb{R}$ means that there exists a constant $C>0$ such that $|f(w)|\leq Cg(w)$ in a neighborhood around 00.
II PROBLEM STATEMENT
Before introducing a favorable problem formulation which will be the basis of the proposed algorithm, a brief overview of related approaches is provided. To this end, the vector of constraint violations for (I) is given by $v(w):=(v_{0},\ldots,v_{N+1})$ where $w$ summarizes all optimization variables in (I) and the stagewise constraint violations are $v_{0}:=x_{1}-\bar{x}_{1}$, $v_{i}:=(x_{i+1}-\phi_{i}(x_{i},u_{i}),[c_{i}(x_{i},u_{i})]^{+})$, $i=1,\ldots,N$, and $v_{N+1}=[c_{N+1}(x_{N+1})]^{+}$.
@@ II PROBLEM STATEMENT
II-A Related Literature
Several successful nonlinear programming solvers implement a procedure to solve the feasibility problem (1). This may be during a feasibility restoration phase which is the case for filterSQP or IPOPT. This phase aims to reduce the infeasibility of the iterates. The feasible iterate solvers FSQP and SEQUOIA solve the feasibility problem before starting their main algorithm from a feasible point. An overview of the feasibility problems that are solved in the different solvers is given in Table I.
TABLE I: Comparison of feasibility problems for different state-of-the-art solvers.
@@ II-A Related Literature
Firstly, consider the feasibility restoration problem of IPOPT as given in Table I. The first observation is that in addition to the minimization of the constraint violation, a regularization term $\frac{\rho}{2}\|w-w^{(0)}\|_{2}^{2}$ with $\rho>0$ is added to the objective. This term aims to find a feasible solution which is close to $w^{(0)}$. If $\rho$ is driven to zero, the solution yields the closest feasible point from $w^{(0)}$. The first term $\|v(w)\|_{1}$ is nonsmooth and is handled in IPOPT or filterSQP by a smooth reformulation introducing slack variables in every constraint. Another key feature of introducing slack variables is that the Linear Independence Constraint Qualification (LICQ) is fulfilled at every feasible point and that the underlying subproblems of the solvers are always feasible. The Hessian of the Lagrangian for the feasibility problem is in general indefinite and it is necessary to properly deal with its negative curvature. An important observation for OCPs is that the gradients of the constraints (I) and (I) are always linearly independent and build a constraint Jacobian with full row rank in an SQP subproblem. Therefore, it is not strictly necessary to introduce additional slack variables for these constraints.
Secondly, looking at the SEQUOIA solver reveals that in the first iteration the function $\frac{1}{2}\|v(w)\|_{2}^{2}$ is minimized to find a feasible point. A favorable property of this formulation is that the function is once continuously differentiable and it has a nonlinear least-squares objective [16]. Due to this structure, a GGN Hessian approximation can be used and no additional reformulation is required to solve the unconstrained optimization problem. Note that this since the objective function is only once continuously differentiable standard analysis needs thus be slightly adapted resulting in the framework of semismooth Newton methods, compare, e.g., [13].
II-B A Favorable Problem Formulation
@@
with $\mathbf{u}:=(u_{0},\ldots,u_{N})$ and $\mathbf{x}:=(x_{1},\ldots,x_{N+1})$. Note that an auxiliary control $u_{0}\in\mathbb{R}^{n_{x}}$ was introduced and the initial condition was softened in the objective function $f_{1}$ such that (II-B) is an unconstrained OCP without initial condition. This choice is motivated by applications such as model predictive control where the solver is often initialized with the solution of the previous time step which is feasible for all constraints except for the initial condition. Another application are Point-to-Point (P2P) motions where the task is to bring a system from initial state $\bar{x}_{1}$ to terminal state $\bar{x}_{N+1}$. Since the initial state and the terminal state should be treated equally, the initial condition constraint is also softened in the objective function.
Eliminating the states and expressing the objective function only with controls is denoted by $F(\mathbf{u})$.
In general, the optimal solution of (II-B) is not unique. All feasible trajectories are an optimal solution to (II-B) and fulfill $f(\mathbf{x},\mathbf{u})=0$.
Formulation (II-B) has the following favorable properties: (i) Keeping the dynamics constraints as constraints preserves the same problem dimension as those of the original OCP and allows LICQ to be satisfied at every feasible point of (II-B); (ii) a Riccati recursion algorithm can efficiently solve the subproblems (II-B). This method has linear complexity with respect to the horizon length [2]; (iii) Due to the nonlinear least-squares objective function in (II-B), a GGN Hessian approximation can be used to approximate the Hessian of the Lagrangian of (II-B); (iv) Since the constraint violation is zero at any optimal solution, the exact Hessian and the GGN Hessian matrices coincide. This implies that local superlinear convergence towards the optimal solution can be achieved.
III FP-DDP ALGORITHM
In this section, the algorithm for finding feasible trajectories, FP-DDP, is described. The proposed method combines DDP with a tailored backtracking line search to guarantee global convergence. In Algorithm 1, the FP-DDP algorithm is presented. The following sections discuss the algorithm’s building blocks.
III-A Differential Dynamic Programming
@@
$A_{i}=\frac{\partial\phi_{i}}{\partial x_{i}}\left(\bar{x}_{i},\bar{u}_{i}\right),~{}B_{i}=\frac{\partial\phi_{i}}{\partial u_{i}}\left(\bar{x}_{i},\bar{u}_{i}\right),$
@@
$\begin{bmatrix}Q_{i}&S_{i}^{\top}\\ S_{i}&R_{i}\end{bmatrix}=H_{i}(\bar{x}_{i},\bar{u}_{i}).$
Defining $R_{0}:=0$, $r_{0}:=0$, $B_{0}:=I$, and $D_{i}:=R_{i}+B_{i}^{\top}P_{i+1}B_{i}$, $d_{i}:=r_{i}+B_{i}^{\top}(P_{i+1}b_{i}+p_{i+1})$, the QP (III-A) can be efficiently solved by a Riccati recursion algorithm, i.e., first a backward sweep calculates the following matrices
@@
$u_{0}$ $=\bar{u}_{0}+\alpha k_{0},$ (6a)
$u_{i}$ $=\bar{u}_{i}+\alpha k_{i}+K_{i}\left(x_{i}-\bar{x}_{i}\right),$ (6b)
$x_{i+1}$ $=A_{i}\left(x_{i}-\bar{x}_{i}\right)+B_{i}\left(u_{i}-\bar{u}_{i}\right)+b_{i}+\bar{x}_{i+1},$ (6c)
for $i=1,\ldots,N$ and $\alpha=1$. Instead of using the linearized dynamics (6c) in the forward rollout, DDP uses the nonlinear dynamics of (II-B) in the forward simulation. This yields the nonlinear rollout
@@
DDP requires initialization at a dynamically feasible point. Due to the nonlinear rollout, all subsequent iterates of DDP are also dynamically feasible. As shown in [14], the optimal objective $f_{\mathrm{QP}}^{*}(\bar{\mathbf{x}},\bar{\mathbf{u}})$ of (III-A) can be efficiently calculated by
$f_{\mathrm{QP}}^{*}(\bar{\mathbf{x}},\bar{\mathbf{u}})\!=\!\!\!\sum_{i=0}^{N}\tfrac{1}{2}k_{i}^{\top}D_{i}k_{i}\!+\!k_{i}^{\top}d_{i}=\!-\!\!\sum_{i=0}^{N}\tfrac{1}{2}d_{i}^{\top}D_{i}^{-1}d_{i}.$ (8)
III-B Regularization
@@
The GGN Hessians for $f_{i},i=1\ldots,n+1$ are denoted by $H_{i}^{\mathrm{GGN}}$. The full GGN Hessian is given by $H^{\mathrm{GGN}}$. The outer convexity of the GGN Hessian is $\tfrac{1}{2}\|[\cdot]^{+}\|_{2}^{2}$ and the inner nonlinearities are $c_{\cdot}(\cdot,\cdot)$. In QP (III-A), the GGN Hessian is combined with an adaptively reduced Levenberg-Marquardt regularization term taken from [3], i.e.,
$H^{(k)}=(H^{\mathrm{GGN}})^{(k)}+\gamma^{(k)}I,$ (9)
@@
where $\gamma^{(k)}:=\mu^{(k)}f(\mathbf{x}^{(k)},\mathbf{u}^{(k)})$ for $\mu^{(k)}>0$ and $I$ denotes the identity matrix. If a full step is taken, the parameter $\mu$ is updated as $\mu^{(k+1)}\leftarrow\max\{\mu_{\min},\bar{\mu}/\lambda\}$ and $\bar{\mu}\leftarrow\mu^{(k)}$, where $\mu_{\min}>0,\lambda>1$ are fixed parameters, $\mu^{(0)}>\mu_{\min}$, and $\bar{\mu}$ is initialized with $\mu^{(0)}$. If a smaller step is taken, then $\mu^{(k+1)}\leftarrow\lambda\mu^{(k)}$. The regularization term vanishes as the iterates approach a feasible point, allowing for superlinear local convergence [3].
If the step size $\alpha$ goes below a given minimum $\alpha_{\min}$ or if the backward Riccati sweep fails, the regularization parameter $\gamma$ is increased similarly to [14], by multiplying $\tau>0$ with $\mu$ and restarting the DDP iteration with the updated regularization parameter.
Since a GGN Hessian approximation is used the proposed algorithm is free of Lagrange multipliers.
III-C Line Search and Step Acceptance

### v2

@@ Abstract
This paper examines the question of finding feasible points to discrete-time optimal control problems. The optimization problem of finding a feasible trajectory is transcribed to an unconstrained optimal control problem. An efficient algorithm, called FP-DDP, is proposed that solves the resulting problem using Differential Dynamic Programming preserving feasibility with respect to the system dynamics in every iteration. Notably, FP-DDP admits global and rapid local convergence properties induced by a combination of a Levenberg-Marquardt method and an Armijo-type line search. An efficient implementation of FP-DDP within acados is compared to established methods such as Direct Multiple Shooting, Direct Single Shooting, and state-of-the-art solvers.
{IEEEkeywords}
Optimal control, optimization algorithms, numerical algorithms
1 Introduction
@@
Reducing infeasibility and finding feasible points is a key building block of constrained optimization algorithms. State-of-the-art nonlinear optimization solvers such as IPOPT [22] or filterSQP [8] employ a feasibility restoration phase to avoid cycling, to overcome infeasible subproblems, and failure of sufficient decrease by iterating towards the feasible set. Feasible iterate solvers, such as SEQUOIA [16] or FSQP [25], require every iterate, including the initial guess, to be feasible. An initial feasibility phase which solves a feasibility problem is often employed to find a feasible initial guess to a generic NLP.
The feasibility-projected Sequential Quadratic Programming (SQP) framework [18], [23] is a feasible iterate method that exploits structure to solve linearly constrained Optimal Control Problems (OCPs). This method employs a trust-region SQP method that uses a feasibility projection to achieve feasibility in every iteration. The feasibility projection is based on Differential Dynamic Programming (DDP) to guarantee feasibility with respect to the dynamics constraints while solving an additional optimization problem to maintain feasibility with respect to linear inequality constraints. Similarly, the method proposed in this paper is based on DDP but relies on reformulating additional inequality constraints as quadratic penalties in the objective function.
@@
DDP, originally introduced in [14], is a method for solving unconstrained OCPs that performs a Riccati backward recursion for a quadratic approximation of the unconstrained OCP and a nonlinear forward simulation to achieve feasibility with respect to the dynamics in every iteration. Local quadratic convergence for exact Hessian DDP was first shown in [15]. The first global convergence proof for line search DDP was presented in [24]. In [2], it was shown that DDP, Direct Multiple Shooting (DMS), and Direct Single Shooting (DSS) recover the same local convergence rate when used with a Generalized Gauss-Newton (GGN) Hessian approximation. Constrained optimization algorithms require a globalization mechanism, where both the decrease of the objective function and the infeasibility of the constraints are controlled. Since the relative weighting of these two terms is nontrivial, DDP is used in this paper as a means of avoiding globalization of the dynamics constraints.
@@
The contributions of this paper are threefold. Firstly, the optimization problem of finding a feasible trajectory is discussed starting from state-of-the-art solvers and a suitable structure-preserving feasibility problem formulation is identified for the special case of OCPs. Secondly, based on the proposed optimization problem, a globally convergent DDP algorithm is introduced. Finally, the novel feasibility algorithm is compared against state-of-the-art methods, DMS, and DSS, on challenging OCPs.
@@
The remainder of this paper is structured as follows. Section 2 discusses the optimization problem of finding a feasible point. In Section 3, the novel feasibility algorithm is introduced. Simulation results are presented in Section 5. Section 6 concludes this paper.
@@
The concatenation of vectors $w_{1}\in\mathbb{R}^{n_{1}}$ to $w_{n}\in\mathbb{R}^{n_{n}}$ is denoted by $(w_{1},\ldots,w_{n})\in\mathbb{R}^{n_{1}+\ldots+n_{n}}$. Iteration indices are denoted by a letter $k$ and superscripts in parentheses, e.g., $w^{(k)}\in\mathbb{R}^{n_{w}}$. For a given iterate $w^{(k)}$, quantities evaluated at that iterate will be denoted with the same superscript, e.g., $f^{(k)}$ or $\nabla f^{(k)}$. In the context of optimal control problems, an OCP stage is denoted by a letter $i$ and subscript, e.g., $x_{i}$ and $u_{i}$. Let $[\cdot]^{+}\colon\mathbb{R}^{n}\to\mathbb{R}^{n},x\mapsto\max\{0,x\}$ denote the componentwise maximum of a vector with the zero vector. For functions $f,g\colon\mathbb{R}^{n_{w}}\to\mathbb{R}$, $f(w)=\mathcal{O}(g(w))$ means that there exists a constant $C>0$ such that $|f(w)|\leq Cg(w)$ in a neighborhood around 00.
2 Problem Statement
@@
2.1 Related Literature
Several successful nonlinear programming solvers solve the feasibility problem (1). This may be during a feasibility restoration phase as for filterSQP and IPOPT. The feasible iterate solvers FSQP and SEQUOIA solve the feasibility problem before starting their main algorithm from a feasible point. An overview of the feasibility problems used in the different solvers is given in Table 1.
Table 1: Comparison of feasibility problems for different state-of-the-art solvers.
@@ 2.1 Related Literature
Firstly, consider the feasibility restoration problem of IPOPT as given in Table 1. The feasibility restoration problem contains in addition to the minimization of the constraint violation, a regularization term $\frac{\rho}{2}\|w-w^{(0)}\|_{2}^{2}$ with $\rho>0$. If $\rho$ is chosen small enough, the IPOPT subproblem is the exact penalty formulation of finding the closest feasible point to $w^{(0)}$ [22]. Note that the correct choice of $\rho$ is not known beforehand. The first term $\|v(w)\|_{1}$ is nonsmooth and is handled in IPOPT or filterSQP by a smooth reformulation introducing slack variables in every constraint. Another key feature of introducing slack variables is that the Linear Independence Constraint Qualification (LICQ) is fulfilled at every feasible point and that the underlying subproblems are always feasible. The Hessian of the Lagrangian for the feasibility problem is in general indefinite and it is necessary to properly deal with its negative curvature. An important observation for OCPs is that the gradients of the constraints (1) are always linearly independent and build a constraint Jacobian with full row rank in an SQP subproblem. Therefore, it is not strictly necessary to introduce slack variables for these constraints.
Secondly, the SEQUOIA solver minimizes $\frac{1}{2}\|v(w)\|_{2}^{2}$ in its first iteration to find a feasible point. A favorable property of this formulation is that it has a nonlinear least-squares objective. Due to this structure, no additional reformulation is required to solve the unconstrained optimization problem and a GGN Hessian approximation can be used. The outer convexity of the GGN Hessian is $\tfrac{1}{2}\|[\cdot]^{+}\|_{2}^{2}$ and the inner nonlinearities are the constraint functions $c_{i},\phi_{i}$. Assuming that the constraints of (1) are continuously differentiable, then the SEQUOIA function is only once continuously differentiable and standard analysis needs to be adapted resulting in the framework of semismooth Newton methods, cf. e.g., [11].
2.2 A Favorable Problem Formulation
@@
with $\mathbf{u}:=(u_{0},\ldots,u_{N})$ and $\mathbf{x}:=(x_{1},\ldots,x_{N+1})$. Note that an auxiliary control $u_{0}\in\mathbb{R}^{n_{x}}$ was introduced and the initial condition was softened in the objective function $f_{1}$ such that (2.2) is an unconstrained OCP without initial condition. This choice is motivated by applications such as model predictive control where the solver is often initialized with the solution of the previous time step which is feasible for all constraints except for the initial condition. Another application are Point-to-Point (P2P) motions where the task is to bring a system from initial state $\bar{x}_{1}$ to terminal state $\bar{x}_{N+1}$. Since the initial state and the terminal state should be treated equally, the initial condition constraint is also softened in the objective function. In general, the optimal solution of (2.2) is not unique. All feasible points of (1) are an optimal solution to (2.2) and fulfill $f(\mathbf{x},\mathbf{u})=0$. If the states are eliminated using (2.2), the objective function only expressed by the controls is denoted by $F(\mathbf{u})$.
Formulation (2.2) has the following favorable properties: (i) keeping the dynamics constraints as constraints preserves the same problem dimension as those of the original OCP and ensures LICQ satisfaction at every feasible point of (2.2); (ii) given the unconstrained OCP structure of (2.2), a Riccati recursion can be used to efficiently solve the subproblems; (iii) due to the nonlinear least-squares objective function in (2.2), a GGN Hessian approximation can be used to approximate the Hessian of the Lagrangian of (2.2); (iv) since the constraint violation is zero at any optimal solution, the exact Hessian and the GGN Hessian coincide. This implies that local superlinear convergence towards the optimal solution can be achieved.
3 FP-DDP Algorithm
In this section, the algorithm for finding feasible trajectories, FP-DDP, is described. The proposed method combines DDP with a tailored backtracking line search to guarantee global convergence. FP-DDP is summarized in Algorithm 1. The following sections discuss its building blocks.
3.1 Differential Dynamic Programming
@@
$A_{i}=\tfrac{\partial\phi_{i}}{\partial x_{i}}\left(\bar{x}_{i},\bar{u}_{i}\right),~{}B_{i}=\tfrac{\partial\phi_{i}}{\partial u_{i}}\left(\bar{x}_{i},\bar{u}_{i}\right),$
@@
$\begin{bmatrix}Q_{i}&S_{i}^{\top}\\ S_{i}&R_{i}\end{bmatrix}=H_{i}(\bar{x}_{i},\bar{u}_{i}).$ (2)
Defining $R_{0}:=0$, $r_{0}:=0$, $B_{0}:=I$, and $D_{i}:=R_{i}+B_{i}^{\top}P_{i+1}B_{i}$, $d_{i}:=r_{i}+B_{i}^{\top}(P_{i+1}b_{i}+p_{i+1})$, the QP (3.1) can be efficiently solved by a Riccati recursion algorithm. First, a backward sweep calculates the following matrices
@@
$u_{0}$ $=\bar{u}_{0}+\alpha k_{0},u_{i}=\bar{u}_{i}+\alpha k_{i}+K_{i}\left(x_{i}-\bar{x}_{i}\right),$ (7a)
$x_{i+1}$ $=A_{i}\left(x_{i}-\bar{x}_{i}\right)+B_{i}\left(u_{i}-\bar{u}_{i}\right)+b_{i}+\bar{x}_{i+1},$ (7b)
for $i=1,\ldots,N$ and $\alpha=1$. Instead of using the linearized dynamics (7b) in the forward rollout, DDP uses the nonlinear dynamics of (2.2) in the forward simulation. This yields the nonlinear rollout
@@
DDP requires $(\bar{\mathbf{x}},\bar{\mathbf{u}})$ to be dynamically feasible. This includes the initial guess. Due to the nonlinear rollout, all subsequent iterates of DDP are also dynamically feasible. As shown in [13], the optimal objective $f_{\mathrm{QP}}^{*}(\bar{\mathbf{x}},\bar{\mathbf{u}})$ of (3.1) can be efficiently calculated by
$f_{\mathrm{QP}}^{*}(\bar{\mathbf{x}},\bar{\mathbf{u}})\!=\!\sum_{i=0}^{N}\tfrac{1}{2}k_{i}^{\top}D_{i}k_{i}\!+\!k_{i}^{\top}\!d_{i}=\!-\!\!\sum_{i=0}^{N}\tfrac{1}{2}d_{i}^{\top}D_{i}^{\shortminus 1}d_{i}.$ (9)
Finally, it is noted that the Riccati recursion is well-defined if $D_{i}$ is invertible. This can be guaranteed by choosing positive definite matrices $H_{i}$ [24]. The subsequent section illustrates how this is ensured by FP-DDP.
3.2 Regularization
@@
The GGN Hessian terms for $f_{i},i=1\ldots,n+1$ are denoted by $H_{i}^{\mathrm{GGN}}$. In (2), the GGN Hessian is combined with an adaptively reduced Levenberg-Marquardt regularization term taken from [3], i.e.,
$H^{(k)}_{i}=(H^{\mathrm{GGN}}_{i})^{(k)}+\gamma^{(k)}I_{i},$ (10)
@@
where $\gamma^{(k)}:=\mu^{(k)}f(\mathbf{x}^{(k)},\mathbf{u}^{(k)})$ for $\mu^{(k)}>0$ and $I_{i}$ denotes the identity matrix of appropriate size. If a full step is taken, the parameter $\mu$ is updated as $\mu^{(k+1)}\leftarrow\max\{\mu_{\min},\tfrac{\bar{\mu}}{\lambda}\}$ and $\bar{\mu}\leftarrow\mu^{(k)}$, where $\mu_{\min}>0,\lambda>1$ are fixed parameters, $\mu^{(0)}>\mu_{\min}$, and $\bar{\mu}$ is initialized with $\mu^{(0)}$. If a smaller step is taken, then $\mu^{(k+1)}\leftarrow\lambda\mu^{(k)}$. The regularization term vanishes as the iterates approach a feasible point, allowing for superlinear local convergence [3]. Since $H^{\mathrm{GGN}}_{i}$ is always positive semi-definite and since the regularization term is positive as long as the algorithm has not converged, $H^{(k)}_{i}$ defined by (10) is ensured to be positive definite. If the step size $\alpha$ goes below a given minimum $\alpha_{\min}$ or if the backward Riccati sweep fails, the regularization parameter $\gamma$ is increased similarly to [13], by multiplying $\lambda$ with $\mu$ and restarting the DDP iteration with the updated regularization parameter. Since a GGN Hessian approximation is used the proposed algorithm is free of Lagrange multipliers.
3.3 Line Search and Step Acceptance